# Boosting Adversarial Transferability with Low-Cost Optimization via Maximin Expected Flatness

### v2

@@ VI-C Evaluation on Defended Models
Figure 4: Parameter ablation study on MEF, evaluating $\gamma$/$\xi$, $\mu_{inner}$/$\mu_{outer}$, and $N$/$T$ impacts on transfer success rates across six challenging targets (Inc-v3 [48], Inc-v4 [57], Xception [61], PiT-S [65], ResMLP [67], Inc-v3adv [37] and Inc-v3ens3 [38]).
VI-D Integrated with Input Augmentation Attacks
We compare MEFH and MEFF against five input augmentation attacks: DI [19], TI [20], SI [22], Admix [25], and SSA [26]. Adversarial examples are generated on Res-50 [9] and Inc-v3 [48], then transferred to six challenging targets: IncRes-v2 [57], Inc-v3ens3 [38], Inc-v4ens4 [38], IncRes-v2ens [38], ViT-L/32 [64], and MLP-Mixer [66]. For fair comparison, each input augmentation attack uses 40 augmented copies per iteration (total $T=10$), requiring 40 backpropagations per step. This makes MEFH (10 iterations, 20 samples) half as costly as input augmentation baselines, while MEFF (20 iterations, 20 samples) matches their computational budget. We further integrate sample transformations from input augmentation methods into MEF variants without introducing additional gradient computations.
Table IV reveals two key findings: (1) MEFH achieves superior transferability (55.12% average success rate) with 50% less computational cost than input augmentation baselines (52.46% for best baseline SSA [26]). When matching the budget, MEFF outperforms SSA by 12.87% (65.33% vs. 52.46%); (2) Combining MEF with input transformations yields substantial gains. For instance, MEFF+TI attains 79.57% average success rate without extra computation, surpassing SSA [26] by 27.11%. This demonstrates that our gradient stabilization framework synergizes effectively with input transformations, particularly benefiting from TI’s translation-invariant patterns. These results validate that MEF provides a more efficient and effective approach than pure input augmentation strategies for enhancing transferability.
TABLE V: Transfer attack success rate (%) comparison across attack paradigms. The best results are bold and the second best results are underlined.
SGM [30] 71.6 35.3 92.8 12.3 12.5 6.1 38.4
LinBP [23] 72.7 47.9 95.4 13.2 13.3 6.8 41.6
ILA [35] 70.5 74.8 97.9 16.5 15.3 9.0 47.3
NAA [24] 82.7 82.4 98.3 30.1 29.6 21.0 57.4
CDA [36] 69.4 83.1 94.5 49.8 52.7 43.3 65.5
MEFH 90.6 88.7 99.7 64.5 63.3 50.4 76.2
MEFF 95.1 94.2 99.9 69.2 65.1 52.2 79.3
VI-E Comparison with Other Types of Attacks
We further validate MEF’s generality by comparing it with three distinct attack paradigms: (1) surrogate refinement methods (SGM [30], LinBP [23]), (2) feature-based approaches (ILA [35], NAA [24]), and (3) generative modeling attack CDA [36]. To ensure fair comparison, all methods generate adversarial examples on Res-50 [9], the prerequisite architecture for LinBP’s skip connection optimization.
As shown in Table V, MEFH achieves 76.2% average success rate across six target models (Inc-v3 [48], Inc-v4 [57], Res-101 [9], Inc-v3ens3 [38], Inc-v3ens4 [38] and IncRes-v2ens [38]), surpassing CDA [36] (65.5%) by 10.7%. Notably, MEFF elevates performance to 79.3%, outperforming CDA [36] (65.5%) by 13.8%. These results demonstrate that our framework consistently outperforms attacks relying on model-specific refinement, intermediate feature alignment, or generative priors.
VI-F Ablation Study
To analyze the impact of key components in MEF, we conduct parametric sensitivity studies on six critical parameters: neighborhood radius $\gamma$, exploration radius $\xi$, outer/inner momentum coefficients ($\mu_{outer}$, $\mu_{inner}$), sampling size $N$, and iteration count $T$. Each parameter’s influence on transferability is evaluated by systematically varying its value while fixing others to default configurations (pre-optimized via exhaustive subset search). Adversarial examples generated on Res-50 are transferred to seven challenging targets spanning standard models (Inc-v3 [48], Inc-v4 [57], Xception [61]), cross-architectures (PiT-S [65], ResMLP [67]), and defended models (Inc-v3adv [37], Inc-v3ens3[38], providing comprehensive insights into parameter robustness across diverse scenarios.
Figure 4 reveals three critical trends. First, both neighborhood radius $\gamma$ and exploration radius $\xi$ (scaled by $\epsilon$) exhibit optimal mid-range values: $\gamma=2\epsilon$ and $\xi=0.15\epsilon$ maximize transferability. Notably, $\xi$ demonstrates higher sensitivity than $\gamma$, yet both optimal values generalize consistently across all targets. Second, while outer momentum $\mu_{outer}=0.5$ provides marginal gains, inner momentum $\mu_{inner}=0.9$ proves indispensable, delivering a 10% success rate boost by balancing historical and current gradient contributions. Third, increasing sampling size $N$ from 5 to 20 improves performance by 6.4%, with diminishing returns beyond $N=15$, while extending iterations $T$ from 5 to 20 yields 37.2% gains before stabilizing. To balance efficiency and performance, we design MEFH (10 iterations, half PGN [8]’s computational cost) and MEFF (20 iterations, matching PGN [8]’s full cost). These results validate that MEF achieves superior transferability through geometry-aware gradient stabilization rather than brute-force sampling or excessive iterations.
VII Conclusion
This work establishes the first theoretical foundation linking multi-order flatness enhancement to adversarial transferability, and proposes a novel flatness-based adversarial attack framework that achieves state-of-the-art transferability performance. In specific, we systematically unify fragmented flatness definitions across existing flatness-enhanced methods, revealing their inconsistent metric adoption and inherent limitations in over-exploration of sensitivity peaks or over-exploitation of local gradients. To resolve these issues, we rigorously formalize the geometric relationship between average-case flatness and transferability gaps, proving that enhancing multi-order average-case flatness minimizes cross-model discrepancies. Based on this theory, we design a maximin expected flatness (MEF) attack that balances worst-neighborhood exploration and local averaging, along with effective sampling methods (NCS) and efficient optimization strategies (GBO) to harmonize gradient exploitation and exploration. Extensive experiments across 22 diverse models and comparisons with 24 state-of-the-art transfer-based attack methods demonstrate our framework’s superior attack success rates and faster computational efficiency than existing approaches. We anticipate this work will inspire broader adoption of flatness principles in AI security and safety research, particularly for enhancing adversarial robustness evaluation and designing next-generation adversarial defense mechanisms.

### v3

@@ VI-E3 Computational Efficiency Analysis
Figure 7: Parameter ablation study on MEF, evaluating $\gamma$$\xi$, $\mu_{inner}$$\mu_{outer}$, and NN/$T$ impacts on transfer success rates across six challenging targets.
VI-E4 Sensitivity to Perturbation Budget
To assess attack robustness under stricter imperceptibility constraints, we evaluate MEF against the top-4 performing baselines (Res-50 $\to$ Inc-v3) while varying $\epsilon$ from $2/255$ to $16/255$. As illustrated in Fig. 8, MEF consistently outperforms competitors across all magnitudes. Notably, at the restrictive budget of $\epsilon=4/255$, MEF achieves a success rate of 41.3%, surpassing the runner-up PGN (31.7%) by a significant margin of 9.6%. This confirms that MEF’s superiority stems from the precise identification of geometrically flat directions rather than reliance on large perturbation norms. For detailed visual inspections across these budgets, please refer to Supp. -J.
Figure 8: Attack success rate under varying perturbation budgets ($\epsilon$).
VI-E5 Parameter Sensitivity
To analyze the impact of key components in MEF, we evaluate six parameters ($\gamma,\xi,\mu_{outer},\mu_{inner},N,T$) across seven diverse targets, ranging from standard CNNs to defended models. The results illustrated in Figure 7 reveals three trends. First, both neighborhood radius $\gamma=2\epsilon$ and exploration radius $\xi=0.15\epsilon$ generalize as optimal values across all targets, demonstrating that MEF is robust to parameter choices and does not require laborious model-specific tuning. Second, inner momentum $\mu_{inner}=0.9$ proves indispensable (delivering a 10% boost), while outer momentum provides marginal gains. Third, increasing sampling size $N$ shows continuous improvement, while extending iterations $T$ to 20 yields substantial gains (37.2%).
TABLE XV: Quantitative Comparison of Zeroth/First-Order Flatness.
Attack Res-50 [9] Inc-v3 [37]
($\overline{R}^{(0)}/\overline{R}^{(1)}$) ($\overline{R}^{(0)}/\overline{R}^{(1)}$)
VMI [1] 29.03 / 0.49 8.32 / 0.64
VNI [1] 23.04 / 0.41 7.74 / 0.59
EMI [4] 25.57 / 0.45 8.77 / 0.62
RAP [7] 15.98 / 0.45 7.80 / 0.62
GNP [5] 50.49 / 0.53 29.37 / 0.86
FEM [3] 12.20 / 0.55 3.52 / 0.79
APP [6] 2.41 / 0.47 1.75 / 0.65
TPA [2] 2.71 / 0.18 1.24 / 0.27
PGN [8] 2.93 / 0.37 2.11 / 0.56
MEFH 2.20 / 0.26 1.09 / 0.27
MEFF 2.34 / 0.10 1.06 / 0.21
VI-E6 Landscape Flatness Analysis
[…]
VII Discussion and Conclusion
Discussion. While MEF significantly outperforms state-of-the-art baselines across challenging cross-task and cross-architecture settings, strictly bridging the structural gap between surrogate and target models remains an open challenge. Moreover, we acknowledge the inherent trade-off between transferability and visual stealthiness in the zero-query setting. Unlike white-box or query-based approaches, transfer-based attacks generally necessitate a larger perturbation budget to ensure the generalization of adversarial features to unknown targets. Nevertheless, MEF establishes a superior trade-off, achieving the most favorable balance between ASR and visual quality among existing transfer-based methods. Addressing these limitations constitutes a meaningful direction for future research. Specifically, future exploration of agnostic priors aims to reconcile the trade-off between cross-domain transferability and visual stealthiness.
Conclusion. This work establishes the first theoretical foundation linking multi-order flatness to adversarial transferability. By unifying fragmented definitions and resolving the exploitation-exploration dilemma, we proposed the Maximin Expected Flatness (MEF) attack. Extensive evaluations demonstrate that MEF achieves state-of-the-art transferability and superior efficiency across diverse models and defenses. We hope this work inspires both deeper and broader adoption of principled flatness optimization in AI security.